(i) State the Existence and Uniqueness theorem for the


differential equation of the first order.


(ii) A home buyer can spend no more than $700 per month on


mortgage payments. Suppose that the interest rate is


7% and that the term of the mortgage is 30 years.


Assume that the interest is compounded continuously


and that payments are also made continuously.


i. Determine the maximum amount that this buyer can


borrow.


ii. Determine the total interest paid during the term


of the mortgage


1
Expert's answer
2021-11-30T16:46:57-0500

(i)

Existence theorem of the 1st order differential equation;

The general first order ODE is"\frac{dy}{dx}=F(x,y)" , "y(x_0)=y_0" .......(i)

Suppose that F (x,y) is a continuous function defined on some region; "R=\{x,y:x_0-\delta<x<x_0+\delta,y_0-\epsilon<y<y_0+\epsilon\}"

containing the point "(x_0,y_0)" . Then there exists a number "\delta_1" (possibly smaller than "\delta") so that a solution "y=f(x)" to equation (i) is defined for "x_0-\delta_1<x<x_0+\delta_1"

Uniqueness theorem of the first order differential equation;

Suppose that both F(x,y) and "\frac{dF}{dy}(x,y)" are continuous functions defined on a region R, "R=\{x,y:x_0-\delta<x<x_0+\delta,y_0-\epsilon<y<y_0+\epsilon\}"

then there exists a number "\delta_2" (possibly smaller than "\delta_1)" so that the solution y=f(x) to equation (i) is the unique solution.


(ii)

(i) m=700

r=7%=0.07

Rate per month ="\frac{0.07}{12}"

n=30yrs=(30*12)months=360months

"M=\frac{P\cdot r(1+r)^n}{(1+r)^n-1}"

"700=\frac{P\cdot (\frac{0.07}{12})(1+\frac{0.07}{12})^{360}}{(1+\frac{0.07}{12})^{360}-1}"

"700=\frac{P\cdot (\frac{0.07}{12})(8.1164975)}{8.1164975-1}"

"700*7.1164975=P(\frac{0.07}{12})(8.1164975)"

0.047346P=4981.5482

P=$105215.81

(ii)I=T-P

T=m*n

=700*360

=252000.00

"\therefore" I=252000.00-105215.81

=$146784.19